# COOL: Efficient and Reliable Chain-Oriented Objective Logic with Neural Networks Feedback Control

### v3

@@ 1 Introduction
Program synthesis is becoming increasingly important in computer science for enhancing development efficiency Gulwani et al. (2017); Jin et al. (2024). Despite the effectiveness of current state-of-the-art methods in dealing with simple tasks, the complexity of modern software demands more advanced and sophisticated approaches Sobania et al. (2022).
@@
To address these challenges, an effective solution must offer programmers fine-grained control and flexible modularity in the synthesis process Groner et al. (2014); Sullivan et al. (2001). First, fine-grained control tailors the synthesis path to specific tasks, ensuring the interpretability of the synthesis process. Secondly, flexible modularity enhances reusability and guarantees the quality of the entire program by ensuring the correctness of the modules Le et al. (2023).
@@
However, these principles are often overlooked in current state-of-the-art program synthesis methods. For example, symbolic approaches such as SyGus Alur et al. (2013), Escher Albarghouthi et al. (2013), and FlashFill++ Cambronero et al. (2023) struggle to scale to complex tasks because their traversal-based Domain-Specific Language (DSL) framework lacks fine-grained control. A compensatory strategy involves using neural networks for guidance or search space pruning, as seen in projects such as Neo Feng et al. (2018), LambdaBeam Shi et al. (2023a), Bustle Odena et al. (2020), DreamCoder Ellis et al. (2023), and Algo Zhang et al. (2023), but the control logic remains disconnected from the programmer. On the other hand, LLM-based projects like CodeGen Nijkamp et al. (2022), CodeX Finnie-Ansley et al. (2022), and Code Llama Roziere et al. (2023) allow programmers to control synthesis through prompt interactions. However, they lack modularity, as all tasks rely on the same LLM, making the logic vulnerable to biases in training data and leading to subtle errors that require manual verification. In summary, there is an urgent need for fine-grained control and flexible modularity to ensure the efficiency and reliability of these methods when tackling complex synthesis tasks.
@@
In this paper, following the principles of fine-grained control and flexible modularity, we present COOL (Chain-Oriented Objective Logic), a neural-symbolic framework for complex program synthesis. At the core of our approach, we introduce the Chain-of-Logic (CoL), which integrates the functions of the activity diagram to enable fine-grained control Gomaa (2011). As illustrated in Figure 1, programmers can precisely organize rules into multiple activities and manage control flow using heuristics and keywords. Additionally, we leverage neural networks on top of CoL to dynamically fine-tune the synthesis process. For this purpose, we introduce Neural Network Feedback Control (NNFC) Turan & Jäschke (2024), which enhances future synthesis by learning from data generated during synthesis and suppresses neural network incorrect predictions through filtering. To ensure modularity, each neural network is bound with a specific CoL DSL, stored in separate library files for clear isolation and easy reuse. Thus, through the combination of CoL and NNFC, COOL achieves high efficiency and reliability when tackling complex synthesis tasks.
@@ 2.2 Neural Network Feedback Control (NNFC)
The quality of the signals generated in the feedback loop directly determines the effectiveness of NNFC. If the error signals are of poor quality, NNFC may not only fail to provide additional improvements but also degrade CoL DSL performance. We ensure the error signal quality through an inner coupling structure within DSNN. As shown in Figure 5 (right), during synthesis tasks, DSNN processes partial programs using a series of sequentially connected neural networks. Each neural network takes both the partial programs and intermediate results from the preceding neural network as input, generating its own predictions. When errors occur in earlier networks, they propagate downstream as noise signals, amplifying at each stage. The difference in the outputs between these neural networks is positively correlated with the accumulated error. To mitigate this, we set a threshold to filter out signals with a significant difference in outputs. Finally, DSNN uses passed signals to generate multi-head outputs (Table 7) to fine-tune the forward flow:
@@ 3 Experiments
We conduct the experiments in two stages to evaluate the improvements introduced by CoL to DSL and to assess how NNFC further enhances performance. First, we carry out static experiments under fixed conditions, including task domain, difficulty level, and neural network. These controlled conditions allow us to accurately measure CoL’s impact on performance. Next, we proceed with dynamic experiments, where conditions vary throughout. This dynamic setup evaluates NNFC’s ability to improve reliability under changing situations.
3.1 Experimental Setup
@@
Benchmarks. We evaluate CoL and NNFC using relational and symbolic tasks with varying difficulty levels, as outlined in Table 1. (Refer to Appendix C for specific task examples.) Specifically, the relational tasks are drawn from the CLUTRR Sinha et al. (2019) dataset, where the goal is to synthesize programs that capture specific target relationships based on human common-sense reasoning. In contrast, the symbolic tasks are generated by GPT Achiam et al. (2023). They involve synthesizing standard quadratic equation programs from non-standard quadratic forms by performing manual calculation steps. Although these tasks are simple for humans, they serve as a straightforward demonstration of how fine-grained control, derived from programmer expertise, can significantly improve program synthesis efficiency.
@@
Metrics. Besides accuracy, we also focus on the following points: (1) CPU Overhead is assessed by the number of tree operations required for synthesis. (2) Memory overhead is assessed by the number of transformation pairs (a partial program paired with the rule to be applied. Appendix A.3).(3) GPU Overhead is measured by the number of neural network invocations. (4) Time overhead is referenced by the actual time spent on program synthesis tasks. (5) Filtering Performance is evaluated by the attenuation ratio of invalid to passed neural network predictions.  [^1] Resource limit: Each partial program must be completed with a cost of at most 1000 transformation pairs and the length of the synthesis path cannot exceed 50.
@@
Chain-of-Logic. We utilize the CoL approach to enhance DSL by making the synthesis process more in line with human problem-solving strategies. For relational tasks, by mirroring the way humans typically reason about family relationships, CoL organizes the synthesis process into stages illustrated in Figure 4. For symbolic tasks, CoL structures the DSL to follow the manual quadratic equation simplification strategy, with stages such as expanding terms, extracting coefficients, permuting terms, and converting equations to standard form. The specific CoL DSL configurations are shown in Table 2, where the significant differences in DSLs highlights the generality of CoL.
Table 2: CoL DSL configurations. The DSL for relational benchmarks has a limited search space and shorter CoL, facing challenges from numerous production rules leading to larger trees. Conversely, the DSL for symbolic benchmarks offers an unlimited search space with a longer CoL, but the many permutation rules increase the risk of cyclic rule applications. (Refer to Appendices I and J for source code snippets of the CoL DSLs used in experiments.)
Benchmark Rules Length of CoL
relational 40 36 2 16 0 4
symbolic 55 17 26 3 11 7
Groups. We use multiple groups to comprehensively evaluate CoL and NNFC (as shown in Table 3). First, in static experiments, we evaluate CoL by comparing DSL groups with and without CoL enhancements. Second, to isolate the impact of heuristic vectors—both as guides and as structuring tools for rule application—we create groups enhanced only by heuristic values. Third, we introduce groups enhanced by neural networks to assess whether combining CoL with neural networks yields better results and to explore the filtering effect of the inner coupling structure. In dynamic
Group Experiment Pretrained DSNN NNFC Inner Coupling Structure
DSL static
✩DSL (Heuristic) static
★CoL DSL static, dynamic
✩DSL+NN static
✩DSL (Heuristic)+NN static
✩ CoL DSL+NN static
✩CoL DSL+NNFC dynamic
✩DSL+NN (Cp) static
✩DSL(Heuristic)+NN (Cp) static
✩CoL DSL+NN (Cp) static
✩CoL DSL+NN (Cp) static
★CoL DSL+NNFC (Cp) dynamic
experiments, we design control groups with and without NNFC to evaluate its impact. Additionally, we include a group without the inner coupling structure to confirm its necessity.
3.2 Static Experiments
@@ 3.3 Dynamic Experiments
Further ablation experiments confirm that reliability provided by NNFC primarily stems from the filtering effect of the inner coupling structure. As shown in Figures 7 and 9, the inner coupling structure reduces the occurrence of accuracy declines due to DSNN mispredictions by 94%. Additionally, the dynamic performance reveals how the inner coupling structure enhances NNFC:
In the scenarios where a DSNN underperforms due to issues such as insufficient training data Mikołajczyk & Grochowski (2018) (as seen in Figure 7, tasks 51-100), inadequate generalization to more challenging tasks Yosinski et al. (2014); Wei et al. (2019) (Figure 7, tasks 301-350), and catastrophic forgetting when tasks from a new domain are learned Kirkpatrick et al. (2017); Van de Ven & Tolias (2019) (Figure 9, tasks 1-100), incorrect predictions lead the actual synthesis path to deviate from the CoL, which in turn causes inefficiency and reduced accuracy. During these phases, for NNFC with the inner coupling structure, the attenuation ratio spikes, indicating that a large percentage of neural network predictions are filtered out. Consequently, the inner coupling structure ensures that the synthesis process adheres to the CoL, effectively mitigating the negative impact of DSNN mispredictions and enhancing reliability.
@@ A.5 A* Search in Program Synthesis
During the exploration phase of program synthesis, we leverage the A* algorithm to perform the heuristic search. This algorithm is renowned for its efficacy in discrete optimization tasks, utilizing heuristic guidance to navigate the search space effectively Hart et al. (1968). Each action or decision is associated with a specific cost in this context. By evaluating the cumulative cost of actions taken so far and the estimated costs of future actions, A* seeks to determine the path with the least overall cost. In our approach, heuristic values promoting forward progression are considered rewards. Therefore, we treat them as negative costs in calculations. Algorithm 1 illustrates the implementation details.
@@
1:procedure A* Search($initialPartialProgram,u_{2}$)
2:     $&openSet\leftarrow\text{priority queue containing only the }\text{initial partial program}\end{aligned}$
3:     $gScore[startPartialProgram]\leftarrow 0$triangleright cost from start
@@
5:     while $openSet\neq\emptyset$ do
6:         $currentProgram\leftarrow openSet.\text{pop}()$triangleright The partial program in openSet with lowest fScore value
7:         if $currentProgram\text{ is complete program}$ then
8:              return Success
9:         end if
10:         for each $neighbor$ of $currentProgram$ do $\triangleright$ Neighbor is a program directly obtained by applying a rule to the current program
[…]
12:              if $tentative_{g}Score<gScore[neighbor]$ then
13:                  $cameFrom[neighbor]\leftarrow current$
14:                  $gScore[neighbor]\leftarrow tentative_{g}Score$
[…]
16:                  if $neighbor\not\in openSet$ then
17:                       $openSet.\text{add}(neighbor)$
18:                  end if
19:              end if
20:         end for
21:     end while
22:     return Failure
23:end procedure
@@ B.2 Neural Network Input
As shown in Figure 5, there are three neural networks in a DSNN. Each network (labeled A, B, and C in their sequential order) takes a partial program as input. The input partial program is represented at the intermediate representation (IR) level in the form of Three-Address Code (TAC) (see Appendix P), allowing program synthesis to be conducted without the constraints of specific DSL syntax or the machine code format of the execution platform Sujeeth et al. (2014). The TAC is then transformed into a graph representation for input to neural networks.
@@ B.4 Neural Network Structure
    Embedding Node Features: We start by employing embedding layers with learning capabilities. These layers convert categorical inputs into dense, continuous vectors, which enhances the stability and efficiency of subsequent processing layers Hrinchuk et al. (2019).
@@
    Graph Feature Extraction: Next, we use a Graph Neural Network (GNN) to extract graph features from each line of TAC code Drori et al. (2022); Wu et al. (2022). To adaptively extract intricate details such as node types, graph attention (GAT) layers are applied after the embedding layers Velickovic et al. (2017).
@@
    Sequential Feature Processing: We adopt Long Short-Term Memory (LSTM) networks to capture the sequential features inherent in TAC Chen et al. (2021); Nye et al. (2020). Recognizing the equal importance of each TAC line, bidirectional LSTM layers are employed following the GAT layers to enrich the contextual understanding Huang et al. (2015).
@@ C.1.3 Multi-Domain Tasks
C.2 Experiment Result
Some experiment results are attached in the appendix without affecting the conclusion.

### v4

@@ Abstract
COOL, Program Synthesis, Machine Learning, Feedback Control
@@ 1 Introduction
Program synthesis is becoming increasingly important in computer science for enhancing development efficiency (Gulwani et al., 2017; Jin et al., 2024). Despite the effectiveness of current state-of-the-art methods in dealing with simple tasks, the complexity of modern software demands more advanced and sophisticated approaches (Sobania et al., 2022).
@@
To address these challenges, an effective solution must offer programmers fine-grained control and flexible modularity in the synthesis process (Groner et al., 2014; Sullivan et al., 2001). First, fine-grained control tailors the synthesis path to specific tasks, ensuring the interpretability of the synthesis process. Secondly, flexible modularity enhances reusability and guarantees the quality of the entire program by ensuring the correctness of the modules (Le et al., 2023).
@@
However, these principles are often overlooked in current state-of-the-art program synthesis methods. For example, symbolic approaches such as SyGus (Alur et al., 2013), Escher (Albarghouthi et al., 2013), and FlashFill++ (Cambronero et al., 2023) struggle to scale to complex tasks because their traversal-based Domain-Specific Language (DSL) framework lacks fine-grained control. A compensatory strategy involves using neural networks for guidance or search space pruning, as seen in projects such as Neo (Feng et al., 2018), LambdaBeam (Shi et al., 2023a), Bustle (Odena et al., 2020), DreamCoder (Ellis et al., 2023), and Algo (Zhang et al., 2023), but the control logic remains disconnected from the programmer. On the other hand, LLM-based projects like CodeGen (Nijkamp et al., 2022), CodeX (Finnie-Ansley et al., 2022), and Code Llama (Roziere et al., 2023) allow programmers to control synthesis through prompt interactions. However, they lack modularity, as all tasks rely on the same LLM, making the logic vulnerable to biases in training data and leading to subtle errors that require manual verification. In summary, there is an urgent need for fine-grained control and flexible modularity to ensure the efficiency and reliability of these methods when tackling complex synthesis tasks.
@@
In this paper, following the principles of fine-grained control and flexible modularity, we present COOL (Chain-Oriented Objective Logic), a neural-symbolic framework for complex program synthesis. At the core of our approach, we introduce the Chain-of-Logic (CoL), which integrates the functions of the activity diagram to enable fine-grained control (Gomaa, 2011). As illustrated in Figure 2, programmers can precisely organize rules into multiple activities and manage control flow using heuristics and keywords. Additionally, we leverage neural networks on top of CoL to dynamically fine-tune the synthesis process. For this purpose, we introduce Neural Network Feedback Control (NNFC) (Turan & Jäschke, 2024), which enhances future synthesis by learning from data generated during synthesis and suppresses neural network incorrect predictions through filtering. To ensure modularity, each neural network is bound with a specific CoL DSL, stored in separate library files for clear isolation and easy reuse. Thus, through the combination of CoL and NNFC, COOL achieves high efficiency and reliability when tackling complex synthesis tasks.
@@ 2.2 Neural Network Feedback Control (NNFC)
The quality of the signals generated in the feedback loop directly determines the effectiveness of NNFC. If the error signals are of poor quality, NNFC may not only fail to provide additional improvements but also degrade CoL DSL performance. We ensure the error signal quality through an inner coupling structure within DSNN. As shown in Figure 5 (right), during synthesis tasks, DSNN processes partial programs using a series of sequentially connected neural networks. Each neural network takes both the partial programs and intermediate results from the preceding neural network as input, generating its own predictions. When errors occur in earlier networks, they propagate downstream as noise signals, amplifying at each stage. The difference in the outputs between these neural networks is positively correlated with the accumulated error. To mitigate this, we set a threshold to filter out signals with a significant difference in outputs. Finally, DSNN uses passed signals to generate multi-head outputs to fine-tune the forward flow.
Figure 6 shows that the DSNN’s multi-head outputs enhance the program synthesis process at all stages (Specific output features in Table 5):
Figure 6: How DSNN’s Multi-Head Outputs Function. When a specific DSL library is loaded, its corresponding DSNN (a specialized network) is loaded to assist in program synthesis by influencing rule application policies for source code partial programs and their derivatives. For instance, if the DSNN bound to the “quadratic" DSL reads a partial program containing a quadratic equation (blue background), it: (1) guides the solver to apply rules specific to its DSL via TDH, (2) prioritizes applying rules on partial programs that are easier to complete using SSPH, and (3) directs the solver to apply specific rules through SGH to progress toward a specified CoL activity. Simultaneously, it prevents the solver from applying quadratic DSL rules to unrelated tasks (e.g., yellow background), ensuring no interference with other DSL synthesis processes. These functions, achieved by modifying heuristic values of DSL-specific rules, demonstrate strong modularity (detailed in Appendix B.6).
@@ 3 Experiments
We conduct the experiments in two stages to evaluate the improvements introduced by CoL to DSL and to assess how NNFC further enhances performance. First, we carry out
@@
static experiments under fixed conditions, including task domain, difficulty level, and neural network. These controlled conditions allow us to accurately measure CoL’s impact on performance. Next, we proceed with dynamic experiments, where conditions vary throughout. This dynamic setup evaluates NNFC’s ability to improve reliability under changing situations.
3.1 Experimental Setup
@@
Benchmarks. We evaluate CoL and NNFC using relational and symbolic tasks with varying difficulty levels, as outlined in Table 1. (Refer to Appendix C for specific task examples.) Specifically, the relational tasks are drawn from the CLUTRR (Sinha et al., 2019) dataset, where the goal is to synthesize programs that capture specific target relationships based on human common-sense reasoning. In contrast, the symbolic tasks are generated by GPT (Achiam et al., 2023). They involve synthesizing standard quadratic equation programs from non-standard quadratic forms by performing manual calculation steps. Although these tasks are simple for humans, they serve as a straightforward demonstration of how fine-grained control, derived from programmer expertise, can significantly improve program synthesis efficiency.
@@
Metrics. (1) Accuracy is the proportion of synthetic tasks that are completed under constraints. [^1] Resource limit: Each partial program must be completed with a cost of at most 1000 transformation pairs and the length of the synthesis path cannot exceed 50. (2) CPU Overhead is assessed by the number of tree operations required for synthesis. (3) Memory overhead is assessed by the number of transformation pairs (a partial program paired with the rule to be applied. Appendix A.3).(4) GPU Overhead is measured by the number of neural network invocations. (5) Time overhead is referenced by the actual time spent on program synthesis tasks. (These metrics measure user performance with DSL libraries, not expert efficiency in writing them.)
@@
Chain-of-Logic. We utilize the CoL approach to enhance DSL by making the synthesis process more in line with human problem-solving strategies. For relational tasks, by mirroring the way humans typically reason about family relationships, CoL organizes the synthesis process into activities illustrated in Figure 4. For symbolic tasks, CoL structures the DSL to follow the manual quadratic equation simplification strategy, with activities such as expanding terms, extracting coefficients, permuting terms, and converting equations to standard form. The specific CoL DSL configurations are shown in Table 7, where the significant differences in DSLs highlights the generality of CoL.
@@
Groups. We use multiple groups to comprehensively evaluate CoL and NNFC (as shown in Table 6). First, in static experiments, we evaluate CoL by comparing DSL groups with and without CoL enhancements. Second, to isolate the impact of heuristic vectors—both as guides and as structuring tools for rule application—we create groups enhanced only by heuristic values. Third, we introduce groups enhanced by neural networks to assess whether combining CoL with neural networks yields better results and to explore the filtering effect of the inner coupling structure. In dynamic experiments, we design control groups with and without NNFC to evaluate its impact. Additionally, we include a group without the inner coupling structure to confirm its necessity.
3.2 Static Experiments
@@ 3.3 Dynamic Experiments
Further ablation experiments confirm that reliability provided by NNFC primarily stems from the filtering effect of the inner coupling structure. As shown in Figures 9 and 10, the inner coupling structure reduces the occurrence of accuracy declines due to DSNN mispredictions by 94%. Additionally, the dynamic performance reveals how the inner coupling structure enhances NNFC:
In the scenarios where a DSNN underperforms due to issues such as insufficient training data (Mikołajczyk & Grochowski, 2018) (as seen in Figure 9, tasks 51-100), inadequate generalization to more challenging tasks (Yosinski et al., 2014; Wei et al., 2019) (Figure 9, tasks 301-350), and catastrophic forgetting when tasks from a new domain are learned (Kirkpatrick et al., 2017; Van de Ven & Tolias, 2019) (Figure 10, tasks 1-100), incorrect predictions lead the actual synthesis path to deviate from the CoL, which in turn causes inefficiency and reduced accuracy. During these phases, for NNFC with the inner coupling structure, the attenuation ratio spikes, indicating that a large percentage of neural network predictions are filtered out. Consequently, the inner coupling structure ensures that the synthesis process adheres to the CoL, effectively mitigating the negative impact of DSNN mispredictions and enhancing reliability.
@@ A.5 A* Search in Program Synthesis
During the exploration phase of program synthesis, we leverage the A* algorithm to perform the heuristic search. This algorithm is renowned for its efficacy in discrete optimization tasks, utilizing heuristic guidance to navigate the search space effectively (Hart et al., 1968). Each action or decision is associated with a specific cost in this context. By evaluating the cumulative cost of actions taken so far and the estimated costs of future actions, A* seeks to determine the path with the least overall cost. In our approach, heuristic values promoting forward progression are considered rewards. Therefore, we treat them as negative costs in calculations. Algorithm 1 illustrates the implementation details.
@@
  function  A* Search ($initialPartialProgram,u_{2}$
@@
     $gScore[startPartialProgram]\leftarrow 0$cost from start}
     $fScore[startPartialProgram]\leftarrow 0$
     while $openSet\neq\emptyset$ do
        $currentProgram\leftarrow openSet.\text{pop( ) }$( ) {Partial program in openSet with lowest fScore value}
        if $currentProgram\text{ is complete program}$ then
           return Success
        end if
        for all $neighbor\in\text{neighbors of }currentProgram$ do
           { Neighbor is obtained by applying a rule to the current program}
           $tentative\_gScore\leftarrow gScore[current]-u_{2}[neighbor]$
           if $tentative\_gScore<gScore[neighbor]$ then
              $cameFrom[neighbor]\leftarrow current$
              $gScore[neighbor]\leftarrow tentative\_gScore$
              $fScore[neighbor]\leftarrow gScore[neighbor]-u_{2}[neighbor]$
              if $neighbor\not\in openSet$ then
                 $openSet.\text{add}(neighbor)$
              end if
           end if
        end for
     end while
     return Failure
  end function
@@ B.2 Neural Network Input
As shown in Figure 5, there are three neural networks in a DSNN. Each network (labeled A, B, and C in their sequential order) takes a partial program as input. The input partial program is represented at the intermediate representation (IR) level in the form of Three-Address Code (TAC) (see Appendix P), allowing program synthesis to be conducted without the constraints of specific DSL syntax or the machine code format of the execution platform (Sujeeth et al., 2014). The TAC is then transformed into a graph representation for input to neural networks.
@@ B.4 Neural Network Structure
    Embedding Node Features: We start by employing embedding layers with learning capabilities. These layers convert categorical inputs into dense, continuous vectors, which enhances the stability and efficiency of subsequent processing layers (Hrinchuk et al., 2019).
@@
    Graph Feature Extraction: Next, we use a Graph Neural Network (GNN) to extract graph features from each line of TAC code (Drori et al., 2022; Wu et al., 2022). To adaptively extract intricate details such as node types, graph attention (GAT) layers are applied after the embedding layers (Velickovic et al., 2017).
@@
    Sequential Feature Processing: We adopt Long Short-Term Memory (LSTM) networks to capture the sequential features inherent in TAC (Chen et al., 2021; Nye et al., 2020). Recognizing the equal importance of each TAC line, bidirectional LSTM layers are employed following the GAT layers to enrich the contextual understanding (Huang et al., 2015).
@@ C.1.3 Multi-Domain Tasks
C.2 Group Configuration
The specific experimental groups of dynamic and static experiments and their respective settings are shown in Table 6:
Group Experiment Trained DSNN NNFC Inner Coupling Structure
DSL static
✩DSL (Heuristic) static
★CoL DSL static, dynamic
✩DSL+NN static
✩DSL (Heuristic)+NN static
✩ CoL DSL+NN static
✩CoL DSL+NNFC dynamic
✩DSL+NN (Cp) static
✩DSL(Heuristic)+NN (Cp) static
✩CoL DSL+NN (Cp) static
✩CoL DSL+NN (Cp) static
★CoL DSL+NNFC (Cp) dynamic
C.3 DSL Configuration
Table 7 shows the CoL DSLs used in the two program synthesis tasks. The CoL length of the DSL in the control group is 1, indicating no activities are created. (Refer to the DSL code snippet in Appendix I and J for activity division and association rules.)
Table 7: CoL DSL configurations. The DSL for relational benchmarks has a limited search space and shorter CoL, facing challenges from numerous production rules leading to larger trees. Conversely, the DSL for symbolic benchmarks offers an unlimited search space with a longer CoL, but the many permutation rules increase the risk of cyclic rule applications.
Benchmark Rules Length of CoL
relational 40 36 2 16 0 4
symbolic 55 17 26 3 11 7
C.4 Experiment Data
Static experiment results are shown in Figure 8:
Singledomain dynamic experiment results are shown in Figure 9:
Multidomain dynamic experiment results are shown in Figure 10: